OpenStudy (anonymous):

Can someone actually explain this to me? x^(2)+5/x+2 + x+1/2x

zepdrix (zepdrix):

Explain what? :) Simplify?

OpenStudy (anonymous):

hi :D

zepdrix (zepdrix):

\[\Large \frac{x^2+5}{x+2}+\frac{x+1}{2x}\]Is that what the problem looks like?

OpenStudy (anonymous):

Just to add it, I forgot the steps... @zepdrix yes that's what is looks like! Hi @sourwing

OpenStudy (mertsj):

\[\frac{x^2+5}{x+2}+\frac{x+1}{2x}\]

OpenStudy (mertsj):

What is the common denominator?

OpenStudy (anonymous):

2, correct?

OpenStudy (mertsj):

The common denominator is 2x(x+2)

OpenStudy (mertsj):

In order to add the fractions, make the denominators the same.

OpenStudy (mertsj):

\[\frac{x^2+5}{x+2}\times \frac{2x}{2x}+\frac{x+1}{2x}\times \frac{x+2}{x+2}\]

OpenStudy (anonymous):

Okay... I will try this. one minute

OpenStudy (anonymous):

2x^(3)+10x/2x^(2)+4 + x^(2)+2/2x^(2)+4x sorry im really bad making the equation on here

OpenStudy (mertsj):

No (x+1)(x+2)=x^2+3x+2

OpenStudy (anonymous):

isn't that the next step?